A Jet Ski rider can travel 16 miles against the current of the lower Mississippi River in 1/2 hour and make the return trip with the current in 1/3 hour. Find the speed of the Jet Ski in still water and the speed of the current.
mph (jet ski in still water)?
mph (current)?

Answer by oberobic(2304)

d = r*t is the basic distance equation: distance = rate * time
.
d = 16 miles
.
r = j - c when going upstream
j -c = jetski speed minus current speed
.
r = j+c when going downstream
.
t = 1/2 hr going upstream
.
16 = 1/2*(j-c)
.
t = 1/3 hr going downstream
.
16 = 1/3(j+c)
.
16=16
so
1/2*(j-c) = 1/3*(j+c)
(j-c)/2 = (j+c)/3
.
Cross multiply
.
3(j-c) = 2(j+c)
3j -3c = 2j + 2c
j = 5c
.
substitute
.
16 = 1/2*(5c-c)
32 = 4c
4c = 32
c = 8
.
Substitute
.
j = 5c
j = 40
.
Check distances traveled in the given times to determine if this answer is correct.
.
1/2*(40-8) = 16
1/3*(40+8) = 16
Correct.
.
Answer: The speed of the jetski in still water is 40 mph, and the speed of the current is 8 mph.